I don't know if this is the right section because I don't know the name of what I am trying to do!! We have to use the numbers 1, 9, 5 and 6 in basic equations and come up with answers from 0-30. For example, (9+1) divided by 5 x 6 = 12. I was wondering if there was an actual formula that I am supposed to be using? Also, can't use decimals, only whole numbers. HELP!!!
Found 2 solutions by solver91311, jim_thompson5910:
Answer by solver91311(12117)



Nope, there is no 'formula' for this. You just have to use your imagination and come up with ways to put those 4 numbers together in thirty-one different ways so that you come up with the results from 0 through 30.





and so on...

Sorry, but you are just going to have to do some thinking. What a concept!

John


Answer by jim_thompson5910(21667)

I don't know of the name either and I'm pretty sure that there isn't a formula. You just have to rearrange the values and try out different things.


Here are all the ones I could find (I couldn't find them all. So I'm not sure if they are all possible or not). I'm sure there are other ways to find them (as they are not unique)


9*(5-6+1)    =   0 
5-1-9+6      =   1 
1*5-9+6      =   2 
1-9+5+6      =   3 
6-(1+9)/5    =   4
15/(9-6)     =   5  ... don't know if this is legal?
9*((5-1)/6)  =   6
9-1+5-6      =   7
1*9+5-6      =   8  
1+9+5-6      =   9
1*9-5+6      =   10  
1+9-5+6      =   11 
((9+1)/5)*6  =   12 ... the example given

5*(9-6)-1    =   14
5*(9-6/1)    =   15
5*(9-6)+1    =   16

9*(6-5+1)    =   18 
9-1+5+6      =   19
1*9+5+6      =   20 
1+9+5+6      =   21  
1-9+5*6      =   22 
6*(9-5)-1    =   23
6*(9-5*1)    =   24
6*(9-5)+1    =   25
5*(6+1)-9    =   26
6*(5+1)-9    =   27
(6+1)*(9-5)  =   28

6*(9-5+1)    =   30


Note: for the question if it was legal or not, I thought about putting the digits 1 and 5 to get 15 (the number fifteen), but I wasn't sure if your teacher allowed it or not.